OpenStudy (anonymous):

Help please, giving it in the form k root 2

OpenStudy (anonymous):

\[2^{\frac{-3}{2}} + 2^{\frac{-1}{2}} +2^{\frac{1}{2}} + 2^{\frac{3}{2}}\]

OpenStudy (anonymous):

@waterineyes

OpenStudy (anonymous):

-2/2 is how much??

OpenStudy (anonymous):

remember: \[\huge \color{green}{x^{\frac{a}{b}} = \sqrt[b]{x^a}}\]

OpenStudy (jiteshmeghwal9):

\[(1/2)^{3/2}+(1/2)^{1/2}+(2)^{1/2}+(2)^{3/2}\]

OpenStudy (anonymous):

What is the first term @zaphod is it -3/2 or -2/2 ??? and last term also..

OpenStudy (anonymous):

-3/2

OpenStudy (anonymous):

Okay...

OpenStudy (anonymous):

last term is 3/2

OpenStudy (anonymous):